A green paint is created by mixing 3 parts of yellow with every 5 parts of blue. How many gallons of yellow paint are needed to make 52 gal of this green paint?

Answer by Theo(13342)

you need 3 parts of yellow for every 5 parts of blue.

you want to make 52 gallons of green paint.

the ratio of yellow to blue is 3:5.

that's a total of 8 parts.

52 / 8 = 6.5

each part requires 6.5 gallons.

3 * 6.5 = 19.5 gallons of yellow.

5 * 6.5 = 32.5 gallons of blue.

19.5 gallons of yellow + 32.5 gallons of blue is equal to 52 gallons of green.

the ratio of yellow to blue is 19.5 / 32.5.

divide both numerator and denominator of this by 6.5 and you get a ratio of 3/5, as you should.